OpenStudy (anonymous):

Which function has x-intercepts of 3 and 8 and contains the point (-1, -2)?

OpenStudy (jdoe0001):

what would be the "roots" of that function you'd think?

OpenStudy (mathstudent55):

Are there choices?

OpenStudy (jdoe0001):

hmm so... got roots? =)

OpenStudy (anonymous):

y=-1/36(x-3)(x-8) y=-1/14(x-3)(x-8) y=-1/18(x-3)(x-8) y=-1/7(x-3)(x-8)

OpenStudy (anonymous):

I don't get what your asking

OpenStudy (jdoe0001):

have you done quadratic equations yet?

OpenStudy (mathstudent55):

A function that has x-intercepts a, b, c, ... has factors of (x - a), (x - b), (x - c), etc. For example, if a function has roots of 2 and -5, then the function is made up of the factors (x - 2) and (x - (-5)), or (x + 5). The function is y = (x - 2)(x + 5). In your case, you have roots 3 and 8. From this you know the function must include these factors: (x - 3) and (x - 8). That means you have to have y = k(x - 3)(x - 8) Now take y = k(x - 3)(x - 8), plug in (-1, -2) and figure out what k is equal to.

OpenStudy (anonymous):

thanks everyone

OpenStudy (mathstudent55):

You're welcome.